If you walk 3/4 mile in 3/5 of an hour, how far do you walk in a full hour
Answer by rfer(16322)

(3/4)/(3/5)=X/1
3/5X=3/4
X=(3/4)/(3/5)
X=3/4*5/3
X=15/12
X=1 1/4 mi